1) Suppose that the demand function of an industry can be described by Q(P) = 1500-

3P. Assume there are two firms with cost function C(g) = 20g.

(a) Determine the equilibrium outcome and profits when firms set quantities sequentially

as in the Stackelberg model. Assume that firm 1 is the leader.

(b) Assume firm 1 is the incumbent in a monopoly market and firm 2 a potential entrant.

Entry implies a fixed cost of F = 300. Can entry be deterred? Will the incumbent firm deter

entry?

(c) Suppose entry implies a fixed cost of F = 30000. Will the incumbent firm deter

entry?

(d) What would firm 1 do if there were no threat of entry? What does this mean for the

incumbent's strategy in part (C)?


1
Expert's answer
2021-08-20T08:49:15-0400
"TR=pQ"

"p=500-\frac{1}{3}Q"

"TR=500Q-\frac{1}{3}Q^2"

"MR=500-\frac{2}{3}Q"

"MC=20"

"500-\frac{2}{3}Q=20"

"Q=720"

"p=260"

"Profit=182,000"

If F=300


"Profit=181,700"

if F-30000


"Profit=152,000"